# Degree-Weighted DeGroot Learning

### v2

@@ 3 General Results
In this section, we examine the impact of the degree-weighted updating rule on learning outcomes—particularly the consensus belief and convergence speed—in the following sequence. We first analyze learning outcomes in expected networks, which are deterministic, before presenting our asymptotic results. While the deterministic case builds on well-established findings in the literature, our primary contribution lies in characterizing the asymptotic properties of learning outcomes in random networks.
3.1 Results for Deterministic and Expected Networks
We begin with necessary definitions. For basic network terminology, we refer readers to standard textbooks such as Jackson et al. (2008).
Definition 1.
A network is called strongly connected if there exists a path connecting any two nodes of the network. A network is aperiodic if there is no integer $k>1$ dividing the length of every cycle of the network.
Definition 2 (Belief Convergence).
Beliefs are said to converge if, for any initial belief profile $b(0)$, the sequence $b(t)$ satisfies
$\lim_{t\to\infty}b_{i}(t)\text{ exists for all }i.$
Definition 3 (Consensus).
The network reaches consensus if beliefs converge to a common value, i.e.,
$\lim_{t\to\infty}b_{i}(t)=b^{*}\quad\forall i.$
where $b^{*}$ is a constant independent of $i$.
Definition 4 (Convergence Speed).
$\max\limits_{\|\mathbf{b}_{0}\|_{2}=1}\|(\mathbf{T}^{t}-\mathbf{T}^{\infty})\mathbf{b}_{0}\|_{2},$ (6)
where $\|\cdot\|_{2}$ denotes the Euclidean norm. This expression quantifies the largest possible deviation between beliefs at time $t$ and the limiting belief, across all initial beliefs whose Euclidean norm is standardized to 1. A larger value indicates slower convergence.
Standard Results in the Literature.
The first well-established result in the literature is that if a network is strongly connected and aperiodic, long-run beliefs not only converge but also reach consensus (see DeGroot, 1974; Meyer, 2000; specifically, Proposition 6 in Golub and Sadler, 2017 for formal proofs and further discussion).
The second standard result in the literature is that the convergence speed of societal beliefs can usually be bound by the second largest eigenvalue in modulus (SLEM) of the learning matrix (see Montenegro, Tetali et al., 2006; Levin and Peres, 2017; see also Proposition 8 in Golub and Sadler, 2017 for a related discussion).
Consensus and Convergence Speed in Degree-Weighted Learning.
[…]
The adaptation of the convergence speed result to our setting is more intricate than in the previous case. Specifically, establishing the inverse relationship between the second largest eigenvalue in modulus (SLEM) and the convergence speed often relies on spectral decomposition, which assumes a diagonalizable learning matrix. A sufficient condition for a real matrix to be diagonalizable is symmetry. However, due to the degree-dependent weighting in our setting, the learning matrix is generally not symmetric. To address this, we show that it is similar to a symmetric matrix, [^14] See Lemma 1 in Appendix B.2.1 allowing us to apply the standard result linking the SLEM to the convergence speed.
We summarize the learning outcomes—particularly the consensus belief and convergence speed for any deterministic network and the expected learning matrix as a special case—in the following proposition.
Proposition 1 (Consensus and Convergence Speed for Deterministic and Expected Degree-Weighted Matrices).
Given any strongly connected and aperiodic adjacency matrix $\mathbf{A}$ with $A_{ij}>0$, [^15] The expected adjacency matrix $\mathbf{R}=\mathbb{E}[\mathbf{A}]$, generated by the stochastic block model, also satisfies this condition. Therefore, in this Proposition, we can replace the matrix $\mathbf{A}$ with $\mathbf{R}$ and $\mathbf{T}$ with $\mathbf{T}^{*}$ as special cases. the degree dependence $\alpha\in\mathbb{R}$, and the corresponding degree-weighted learning matrix $\mathbf{T}=\mathbf{T}(\mathbf{A};\alpha)$ as defined in Equation (3), we have:
@@ 3.1 Results for Deterministic and Expected Networks
    $s_{i}=\frac{\sum\limits_{j}A_{ij}\,d_{j}^{\alpha}(\mathbf{A})\,d_{i}^{\alpha}(\mathbf{A})}{\sum\limits_{i^{\prime},j}A_{i^{\prime}j}\,d_{j}^{\alpha}(\mathbf{A})\,d_{i^{\prime}}^{\alpha}(\mathbf{A})}.$ (7)
@@
    Let $|\lambda_{2}(\mathbf{T})|$ denote the second largest eigenvalue (in modulus) of $\mathbf{T}$ and let $\|\cdot\|_{2}$ denote the Euclidean norm. Then, at any time $t$,
    $\max\limits_{\|\mathbf{b}_{0}\|_{2}=1}\|(\mathbf{T}^{t}-\mathbf{T}^{\infty})\mathbf{b}_{0}\|_{2}=|\lambda_{2}(\mathbf{T})|^{t}.$ (8)
@@
Proof.
See Appendix B.3. ∎
We note a few things here. Our first result, which characterizes agent influence with degree-dependent weighting, helps us leverage previous results established in the literature that characterize conditions under which the consensus belief coincides with the true state (e.g., Lemma 1, p. 124 in Golub and Jackson, 2010). We devote Section 4.1 to a discussion of a wisdom-of-crowds type result.
Secondly, from the characterization of influence provided in Equation (7), one can see that when $\alpha=0$, we return to the equal-weighting case, and the agent’s influence is exactly as delivered in the classical literature (such as Theorem 6, p. 935 in DeMarzo, Vayanos and Zwiebel, 2003 and p. 122 in Golub and Jackson, 2010). Another special case is $\alpha=-1$. A brief calculation shows that, in this case, agents’ influence remains unequal; that is, their impact through degree is not fully offset by the popularity-resistant learning rule $\alpha=-1$ (see Section 4.1 for further discussion).
Lastly, by demonstrating that the second largest eigenvalue in modulus (SLEM) remains inversely related to the convergence speed in our case of degree-dependent weighting, we justify the examination of the asymptotic properties of the SLEM of a random learning matrix $\mathbf{T}$ and the comparative statics of $\alpha$ on the SLEM of the expected learning matrix $\mathbf{T^{*}}$.
3.2 Main Result: Random Networks and Asymptotic Results
In the previous section, we discussed standard results in the literature on consensus and the speed of belief convergence for deterministic networks. These results were extended to the expected adjacency and learning matrices in our case with degree-dependent weighting, $\mathbf{R}=\mathbb{E}[\mathbf{A}]$ and $\mathbf{T}^{*}=\mathbf{T}(\mathbf{R},\alpha)$, given any degree-dependence $\alpha$.
In this section, we focus on random networks and establish our key asymptotic results. In particular, we demonstrate that the learning outcomes corresponding to the realized networks, $\mathbf{A}$ and $\mathbf{T}$, are, in a probabilistic sense, sufficiently close to those corresponding to the expected networks, $\mathbf{R}$ and $\mathbf{T}^{*}$. [^16] In Appendix D, we show that our main results are also robust to further (minor) perturbations to the adjacency matrix $\mathbf{A}$.
Theorem 1.
Given any degree-dependence parameter $\alpha\in\mathbb{R}$ and a random network $\mathbf{A}(\mathbf{P},\mathbf{n})$, there exists a positive constant $\tilde{C}$, independent of $n$, such that for all $n>0$, the following results hold for the realized learning matrices $\mathbf{T}(\mathbf{A},\alpha)$ and the expected learning matrix $\mathbf{T}^{*}=\mathbf{T}(\mathbf{R},\alpha)$:
@@ 3.2 Main Result: Random Networks and Asymptotic Results
    For influences,
    $\operatorname{\mathbb{P}}\Bigl{\{}\|\mathbf{s}(\mathbf{T})-\mathbf{s}(\mathbf{T}^{*})\|\geq\tilde{C}\sqrt{n\log n}\,(\min_{i}d_{i}(\mathbf{R}))^{-1}\Bigr{\}}\leq\frac{16}{n^{2}}\,.$
@@
The formal proof is involved and is provided in Appendix B.4. Since Theorem 1 constitutes our primary technical contribution, we briefly outline the key ideas of the proof and indicate how our approach diverges from the existing literature.
The key steps in proving our asymptotic result are as follows:
  1. 1.
    To establish that $\mathbf{T}$ and $\mathbf{T}^{*}$ are similar to symmetric matrices, where
    $\mathbf{T}=\mathbf{D}_{1,\mathbf{A}}^{-1}\mathbf{A}\mathbf{D}_{2,\mathbf{A}},\quad\mathbf{T}^{*}=\mathbf{D}_{1,\mathbf{R}}^{-1}\mathbf{R}\mathbf{D}_{2,\mathbf{R}},$
    and $\mathbf{D}_{i,\mathbf{A}}$ and $\mathbf{D}_{i,\mathbf{R}}$ are diagonal matrices. [^17] See Lemma 1 in Appendix B.4, and Equations (15) and (18) in the Appendix for the above expressions.
  2. 2.
    To bound $|{\lambda_{2}}(\mathbf{T})-{\lambda_{2}}(\mathbf{T}^{*})|$, we apply Weyl’s inequality, which relies on the preceding result. This allows us to reduce the problem to establishing bounds for $\|\mathbf{A}-\mathbf{R}\|$, $\|\mathbf{D}_{1,\mathbf{A}}-\mathbf{D}_{1,\mathbf{R}}\|$, and $\|\mathbf{D}_{2,\mathbf{A}}-\mathbf{D}_{2,\mathbf{R}}\|$.
  3. 3.
    Bounding $\|\mathbf{A}-\mathbf{R}\|$ is standard in the literature (Golub and Jackson, 2012; Dasaratha, 2020; Mostagir and Siderius, 2021).
  4. 4.
    The difficulty lies in bounding $\|\mathbf{D}_{1,\mathbf{A}}-\mathbf{D}_{1,\mathbf{R}}\|$ and $\|\mathbf{D}_{2,\mathbf{A}}-\mathbf{D}_{2,\mathbf{R}}\|$, which are nonlinear in $\mathbf{A}$ and $\mathbf{R}$ in our degree-weighted framework.
  5. 5.
    Our approach: We use a leave-one-out technique (e.g., Abbe et al., 2020), replacing $d_{j}^{\alpha}(\mathbf{A})$ with $d_{j,-i}^{\alpha}(\mathbf{A})$ so that $A_{ij}$ is independent of $d_{j,-i}^{\alpha}(\mathbf{A})$, yielding $d_{j,-i}^{\alpha}(\mathbf{A})\approx d_{j}^{\alpha}(\mathbf{A})$ and thus
    $(\mathbf{D}_{1,\mathbf{A}})_{ii}\approx\sum_{j}A_{ij}\,d_{j,-i}^{\alpha}(\mathbf{A}).$
    Noting that, for fixed $i$, the collection $\{A_{ij}:j\neq i\}$ is independent of $\{d_{j,-i}(\mathbf{A}):j\neq i\}$, it follows that, conditional on $\{d_{j,-i}(\mathbf{A}):j\neq i\}$, the sum $\sum_{j\neq i}A_{ij}\,d_{j,-i}^{\alpha}(\mathbf{A})$ consists of independent random variables.
    This observation enables us to control the convergence rate of $(\mathbf{D}_{1,\mathbf{A}})_{ii}$ using standard concentration inequalities for sums of independent random variables. [^18] We refer the reader to Appendix B.4 for the complete proof.
Remark 1.
@@
$T_{ij}=\frac{A_{ij}d_{j}^{\alpha_{j}}}{\sum\limits_{j^{\prime}}A_{ij^{\prime}}d_{j^{\prime}}^{\alpha_{j^{\prime}}}},\quad T^{*}_{ij}=\frac{R_{ij}d_{j}^{\alpha_{j}}(\mathbf{R})}{\sum\limits_{j^{\prime}}R_{ij^{\prime}}d_{j^{\prime}}^{\alpha_{j^{\prime}}}(\mathbf{R})},\quad s_{i}=\frac{\sum\limits_{j}R_{ij}\,d_{j}^{\alpha_{j}}(\mathbf{R})d_{i}^{\alpha_{j}}(\mathbf{R})}{\sum\limits_{i^{\prime},j}R_{i^{\prime}j}d_{j}^{\alpha_{j}}(\mathbf{R})d_{i^{\prime}}^{\alpha_{j}}(\mathbf{R})},\quad\forall i,j,$
where $\max_{i,j}|\alpha_{i}-\alpha_{j}|\leq c$ for some positive constant $c$.
Proof.
See Appendix B.5. ∎
Theorem 1 lays the foundation for further discussion of economic insights in the degree‐weighted learning framework. In particular, to understand key learning outcomes, we are able to focus on the expected matrices rather than all possible realized networks. Together with Proposition 1 in the previous section, we know that influences (and later, the wisdom of a society) are captured by Equation (7), and that the convergence speed is inversely related to $|\lambda_{2}|$ of the expected learning matrix $\mathbf{T}^{*}$, as defined in Equation (8). In the next section, we provide a series of examples to highlight the impacts of degree‐weighted learning, as captured by $\alpha$, on societal wisdom, convergence speed, and a potential tradeoff between the two.
@@ 4 Applications
In this section, we apply the asymptotic results established above to three questions. First, we examine how the degree-dependent learning rule affects societal wisdom and find that its impact is largely monotonic. Next, we study its effect on the speed of learning, which displays a non-monotonic pattern. Lastly, we discuss a potential trade-off between societal wisdom and learning speed.
4.1 Influences and Wisdom of the Crowd
This section explores the concepts of influence and wisdom. [^19] We thank the editor and the anonymous referees for suggesting the discussion on wisdom. We begin by considering special cases of the degree-dependence parameter $\alpha$ to illustrate its impact on agents’ influences. Adopting the notion of wisdom introduced by Golub and Jackson (2010), we then show that variations in $\alpha$ are powerful enough to shift a society from wisdom to non-wisdom, or vice versa. Moreover, this effect is monotonic: increasing $\alpha$ toward a popularity-favoring learning rule leads to non-wisdom, while decreasing $\alpha$ toward a popularity-resistant rule promotes wisdom.
Influences With Special Cases of $\alpha$.
Recall that in Equation (7) of Proposition 1, we obtain the following characterization of the impact of $\alpha$ on influences for any deterministic network $\mathbf{A}$ that satisfies the imposed regularity conditions:
$s_{i}=\frac{\sum\limits_{j}A_{ij}\,d_{j}^{\alpha}d_{i}^{\alpha}}{\sum\limits_{i^{\prime},j}A_{i^{\prime}j}d_{j}^{\alpha}d_{i^{\prime}}^{\alpha}}.$
A notable special case is $\alpha=0$ (equal weighting). As shown in the literature (e.g., DeMarzo, Vayanos and Zwiebel, 2003 and Golub and Jackson, 2010), we have
$s_{i}=\frac{d_{i}}{\sum_{j}d_{j}},$
so that the influence of agent $i$ is proportional to his or her degree.

### v3

@@ 3 General Results
In this section, we present our main theoretical result for the degree-weighted updating rule: an asymptotic theorem showing that, as the network grows large, the learning outcomes—specifically, the consensus belief and the convergence speed of the realized learning matrix $\mathbf{T}(\mathbf{A};\alpha)$—converge to those of the expected learning matrix $\mathbf{T}^{*}(\mathbf{R};\alpha)$. This result allows us to focus on the expected case, $\mathbf{T}^{*}(\mathbf{R};\alpha)$, when discussing the comparative statics of learning outcomes in Section 4.
As preparation, we begin by characterizing consensus beliefs and convergence speed for deterministic networks, with the expected network as a special case. The first part of Proposition 1 adapts the standard characterization of consensus beliefs in DeGroot learning to our degree-weighted setting, and the second part generalizes Lemma 2 of Golub and Jackson (2012).
Proposition 1 (Learning Outcomes for Deterministic and Expected Matrices)
Given any primitive matrix $\mathbf{A}$ such that $\min_{i}d_{i}(\mathbf{A})>0$, [^16] The expected adjacency matrix $\mathbf{R}={\rm I}\kern-1.79993pt{\rm E}\mathbf{A}$, generated by the stochastic block model, is primitive under Assumptions 13 (see Proposition A.1 in the Appendix). Therefore, in this proposition we can replace $\mathbf{A}$ with $\mathbf{R}$ and $\mathbf{T}$ with $\mathbf{T}^{*}$. Assumptions 13 also imply that $\mathbf{A}$ is primitive with probability tending to one, which in turn implies that the precondition for Proposition 1 holds with high probability. we have:
@@
    (5) $s_{i}=\frac{\sum\limits_{j}A_{ij}\,d_{j}^{\alpha}(\mathbf{A})\,d_{i}^{\alpha}(\mathbf{A})}{\sum\limits_{i^{\prime},j}A_{i^{\prime}j}\,d_{j}^{\alpha}(\mathbf{A})\,d_{i^{\prime}}^{\alpha}(\mathbf{A})}.$
@@
    Let $|\lambda_{2}(\mathbf{T})|$ denote the second-largest eigenvalue (in modulus) of $\mathbf{T}$ and let $\|\mathbf{v}\|_{\mathbf{s}}=\sqrt{\sum_{i=1}^{n}s_{i}\mathbf{v}_{i}^{2}}$, where $\mathbf{v}_{i}$ is the $i$-th entry of $\mathbf{v}$ and $s_{i}$ is agent $i$’s influence defined in Equation (5). Then, at any time $t$,
    (6) $\max\limits_{\|\mathbf{b}_{0}\|_{\mathbf{s}}=1}\|(\mathbf{T}^{t}-\mathbf{T}^{\infty})\mathbf{b}_{0}\|_{\mathbf{s}}=|\lambda_{2}(\mathbf{T})|^{t}.$
@@
We note a few things here. It is a standard result in the literature that if a network is primitive, long-run beliefs converge to consensus, [^17] See DeGroot (1974) and Meyer (2000); specifically, Proposition 6 in Golub and Sadler (2016). and that the convergence speed of beliefs is determined by the second-largest eigenvalue in modulus (SLEM) of the learning matrix. [^18] See Montenegro and Tetali (2006) and Levin and Peres (2017); see also Proposition 8 in Golub and Sadler (2016). Our characterization of agents’ influences and convergence speed in the degree-weighted deterministic setting (Proposition 1) provides the foundation for the asymptotic result below and for applying our framework to a wisdom-of-crowds result and the comparative statics in Section 4.
We are now ready to introduce our key technical result. We show that the learning outcomes corresponding to the realized networks, $\mathbf{A}$ and $\mathbf{T}$, are, in a probabilistic sense, sufficiently close to those corresponding to the expected networks, $\mathbf{R}$ and $\mathbf{T}^{*}$. [^19] In Online Appendix O3, we show that our main results are robust to minor perturbations of the adjacency matrix $\mathbf{A}$.
Theorem 1
Under Assumptions 13, for any degree dependence $\alpha\in\mathbb{R}$ and any random network $\mathbf{A}(\mathbf{P},\mathbf{n})$, there exists a positive constant $\tilde{C}$, independent of $n$, such that for all $n>0$, the following results hold for the realized learning matrix $\mathbf{T}(\mathbf{A},\alpha)$ and the expected learning matrix $\mathbf{T}^{*}=\mathbf{T}(\mathbf{R},\alpha)$.
@@
    For influences, [^20] Recall that $\tau_{n}=\min_{i}{d_{i}(\mathbf{R})}/{n}$. Assumption 1 implies that ${\sqrt{\log n}}/{(\tau_{n}\sqrt{n})}$ goes to 0 as $n\rightarrow\infty$.
    $\operatorname{\mathbb{P}}\Bigl\{\|\mathbf{s}(\mathbf{T})-\mathbf{s}(\mathbf{T}^{*})\|_{2}\geq\tilde{C}\frac{\sqrt{\log n}}{\tau_{n}\sqrt{n}}\Bigr\}\leq\frac{16}{n^{2}}\,.$
@@
 [^21] The formal proof is involved and is provided in Appendix 4. We briefly outline the key ideas here and indicate how our approach diverges from the existing literature.
The critical step of the proof is to show that the symmetry of the adjacency matrix $\mathbf{A}$ ensures that the learning matrices $\mathbf{T}$ and $\mathbf{T}^{*}$ are similar to symmetric matrices. This property is pivotal, as it enables the application of eigenvalue perturbation methods specifically designed for symmetric or symmetrically similar matrices.
Specifically, we express $\mathbf{T}=\mathbf{D}_{1,\mathbf{A}}^{-1}\mathbf{A}\mathbf{D}_{2,\mathbf{A}},$ and $\mathbf{T}^{*}=\mathbf{D}_{1,\mathbf{R}}^{-1}\mathbf{R}\mathbf{D}_{2,\mathbf{R}}$ where $\mathbf{D}_{i,\mathbf{A}}$ and $\mathbf{D}_{i,\mathbf{R}}$ are diagonal matrices that depend nonlinearly on degrees. Lemma B.1 in the appendix shows how the symmetry of the adjacency matrix $\mathbf{A}$ implies that $\mathbf{T}$ and $\mathbf{T}^{*}$ are similar to symmetric matrices.
With symmetry established, we then bound the difference $|\lambda_{2}(\mathbf{T})-\lambda_{2}(\mathbf{T}^{*})|$ through Weyl’s inequality, a standard perturbation result that requires symmetry (or symmetric similarity) of the involved matrices. This reduces our analysis to deriving suitable bounds on the matrix differences $\|\mathbf{A}-\mathbf{R}\|$, $\|\mathbf{D}_{1,\mathbf{A}}-\mathbf{D}_{1,\mathbf{R}}\|$, and $\|\mathbf{D}_{2,\mathbf{A}}-\mathbf{D}_{2,\mathbf{R}}\|$. [^22] For a matrix $\mathbf{B}$, $\|\mathbf{B}\|=\sup_{\|\mathbf{v}\|=1}=\|\mathbf{B}\mathbf{v}\|$, where $\|\mathbf{v}\|$ is Euclidean norm of vector $\mathbf{v}$.
Bounding $\|\mathbf{A}-\mathbf{R}\|$ is relatively straightforward and follows standard procedures from the literature (e.g., Golub and Jackson, 2012; Dasaratha, 2020; Mostagir and Siderius, 2021). However, the nonlinearity introduced by degree weighting presents significant difficulties in bounding the diagonal matrices $\mathbf{D}_{1,\mathbf{A}}$ and $\mathbf{D}_{2,\mathbf{A}}$.
To handle this nonlinearity, we employ a leave-one-out strategy, inspired by techniques such as those developed by Abbe et al. (2020). This approach involves replacing $d_{j}^{\alpha}(\mathbf{A})$ with $d_{j,-i}^{\alpha}(\mathbf{A})$, which excludes the direct influence of the element $A_{ij}$. Consequently, the modified degree $d_{j,-i}^{\alpha}(\mathbf{A})$ becomes independent of $A_{ij}$, allowing us to approximate the diagonal entries as $(\mathbf{D}_{1,\mathbf{A}})_{ii}\approx\sum_{j}A_{ij}\,d_{j,-i}^{\alpha}(\mathbf{A}).$
Given this construction, the sum $\sum_{j\neq i}A_{ij}\,d_{j,-i}^{\alpha}(\mathbf{A})$ becomes a sum of independent random variables, conditional on $\{d_{j,-i}(\mathbf{A}):j\neq i\}$. This step allows us to employ standard concentration inequalities such as Bernstein inequality (Vershynin, 2018) to control the convergence rates of the diagonal entries of $\mathbf{D}_{1,\mathbf{A}}$ and $\mathbf{D}_{2,\mathbf{A}}$. This concludes the key steps of the proof. [^23] We briefly outline the intuition for each assumption in the context of this theorem. The proof relies on bounding the deviation between the eigenvectors of the random matrix $\mathbf{T}$ and the expected matrix $\mathbf{T}^{*}$. This bound generally takes the form $\text{Error}\lesssim{\text{Noise}}/{\text{Gap}}$. Assumptions 1 and 2 control the noise (numerator): they ensure sufficiently dense connections so that the random matrix concentrates tightly around its mean via the Law of Large Numbers. Assumption 3 guarantees the gap (denominator): it ensures the spectral gap of the expected matrix remains strictly bounded away from zero, thereby providing the structural stability required for the eigenvectors to be robust to noise. $\Box$
Remark 1
@@
$T_{ij}=\frac{A_{ij}d_{j}^{\alpha_{j}}}{\sum\limits_{j^{\prime}}A_{ij^{\prime}}d_{j^{\prime}}^{\alpha_{j^{\prime}}}},\quad T^{*}_{ij}=\frac{R_{ij}d_{j}^{\alpha_{j}}(\mathbf{R})}{\sum\limits_{j^{\prime}}R_{ij^{\prime}}d_{j^{\prime}}^{\alpha_{j^{\prime}}}(\mathbf{R})},\quad s_{i}=\frac{\sum\limits_{j}R_{ij}\,d_{j}^{\alpha_{j}}(\mathbf{R})d_{i}^{\alpha_{i}}(\mathbf{R})}{\sum\limits_{i^{\prime},j}R_{i^{\prime}j}d_{j}^{\alpha_{j}}(\mathbf{R})d_{i^{\prime}}^{\alpha_{i^{\prime}}}(\mathbf{R})},\quad\forall i,j,$
where $\max_{i,j}|\alpha_{i}-\alpha_{j}|\leq c$ for some positive constant cc. [^24] See Appendix 4.2 for a proof.
Theorem 1 lays the foundation for further discussion of economic insights in the degree‐weighted learning framework. In particular, to understand key learning outcomes, we are able to focus on the expected matrices rather than all possible realized networks. Together with Proposition 1 in the previous section, we know that influences (and later, the wisdom) are captured by Equation (5), and that the convergence speed is inversely related to $|\lambda_{2}|$ of the expected learning matrix $\mathbf{T}^{*}$, as defined in Equation (6). In the next section, we provide a series of examples to highlight the impacts of degree‐weighted learning, as captured by $\alpha$, on societal wisdom, convergence speed, and a potential trade-off between the two.
@@ 4 Applications
In this section, we apply the asymptotic results established above to three questions. First, we examine how the degree-dependent learning rule affects societal wisdom and find that its impact is largely monotonic. Next, we study its effect on the speed of learning, which displays a non-monotonic pattern. Lastly, we discuss a potential trade-off between societal wisdom and learning speed. With Theorem 1, we can focus on the comparative statics of expected networks rather than of specific realized networks.
A Two-Type Model
For analytical tractability, we consider a class of stochastic block models with two levels of expected degree. Specifically, suppose there are $m$ groups with sizes $n_{1}\neq n_{2}=\cdots=n_{m}$, where within-group linking probabilities are given by $p_{kk}=p$ and across-group probabilities by $p_{kl}=q$ for $l\neq k$, with $p\neq q$. Agents therefore fall into two types: group 1 has expected degree $d_{1}^{*}=n_{1}p+(m-1)n_{2}q$, and each group $k>1$ has expected degree $d_{2}^{*}=n_{1}q+n_{2}p+(m-2)n_{2}q$. We refer to the group(s) with larger expected degree as the high-degree group(s), and those with smaller expected degree as the low-degree group(s). [^25] Group 1 is the high-degree group whenever $d_{1}^{*}-d_{2}^{*}=(n_{1}-n_{2})(p-q)>0$.
This structure, though simple, accommodates three cases distinguished by the numbers of high-degree ($m_{h}$) and low-degree ($m_{l}$) groups: (i) $m_{h}=1$ and $m_{l}=1$; (ii) $m_{h}>1$ and $m_{l}=1$, capturing situations in which high-degree agents represent different campaigns or hold conflicting views; and (iii) $m_{h}=1$ and $m_{l}>1$, corresponding to environments with multiple low-degree groups that may reflect cultural, ethnic, or ideological divisions.
4.1 Influence and Wisdom of the Crowd
This section discusses influence and wisdom. [^26] We thank the editor and the anonymous referees for suggesting this discussion. Our goal is to combine insights from Golub and Jackson (2010) with our main technical result, Theorem 1, to characterize the effect of degree dependence on societal wisdom.
Specifically, we consider a sequence of realized networks generated by a stochastic block model, and analyze the corresponding sequence of learning matrices. The wisdom criterion, established in Golub and Jackson (2010), states that society is wise if and only if $\lim\limits_{n\rightarrow\infty}\max\limits_{i\leq n}s_{i}(\mathbf{T}(n))=0$, where $s_{i}(\mathbf{T}(n))$ is agent $i$’s influence under the learning matrix $\mathbf{T}(n)$. Our Theorem 1 shows that, under Assumptions 13, the influence vector of the realized network converges to that of the expected network as $n\rightarrow\infty$. Therefore, if $\lim\limits_{n\rightarrow\infty}\max\limits_{i\leq n}s_{i}(\mathbf{T}^{*}(n))=0$ for the expected matrix $\mathbf{T}^{*}(n)$, then, with high probability, the realized sequence of networks will also be wise. Conversely, if $\liminf\limits_{n}\max\limits_{i\leq n}s_{i}(\mathbf{T}^{*}(n))>0$, wisdom fails with high probability. This logic ensures that our comparative static results with respect to $\alpha$ hold with high probability in large networks.
Influences With Special Cases of $\alpha$
We begin by considering special cases of the degree-dependence parameter $\alpha$ to illustrate its effect on agents’ influence. Recall that in Equation (5) of Proposition 1, we obtain the following characterization of the impact of $\alpha$ on influence for any deterministic matrix $\mathbf{A}$ that satisfies the imposed conditions: $s_{i}={(\sum_{j}A_{ij}\,d_{j}^{\alpha}d_{i}^{\alpha})}/{(\sum_{i^{\prime},j}A_{i^{\prime}j}\,d_{j}^{\alpha}d_{i^{\prime}}^{\alpha})}.$
A notable special case is $\alpha=0$ (equal weighting). As shown in the literature (e.g., DeMarzo, Vayanos and Zwiebel, 2003 and Golub and Jackson, 2010), we have $s_{i}={d_{i}}/{\sum_{j}d_{j}}$, so that the influence of agent $i$ is proportional to their degree.